# Adaptive debiased SGD in high-dimensional GLMs with steaming data

### v1

@@ 2.4 Variance and confidence interval
To construct confidence interval, we still need to find the variance of ${\widehat{\bm{\beta}}}^{(k)}_{j,\text{de}}$. To this end, we design another scalar-valued summary statistic $a_{5}^{(k)}=\sum_{i=n_{l}+1}^{k}(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})-y_{i}\right\}^{2}$. Then the estimated standard error of the ADL estimator is computed as
$\widehat{\nu}^{(k)}_{j}=\sqrt{a_{5}^{(k)}}/|a_{4}^{(k)}|.$ (15)
Consequently, based on (13) and (15), the $(1-\alpha)$-confidence interval for $\bm{\beta}_{j}^{*}$ with $10<\alpha<$ 1 at time $k$ is given by:
$\left({\widehat{\bm{\beta}}}^{(k)}_{j,\text{de}}-z_{\alpha/2}\widehat{\nu}^{(k)}_{j},{\widehat{\bm{\beta}}}^{(k)}_{j,\text{de}}+z_{\alpha/2}\widehat{\nu}^{(k)}_{j}\right),\ k>n_{l},$ (16)
where $z_{\alpha/2}$ is the $\{1-(\alpha/2)\}$-quantile of the standard normal distribution. We present the pseudocode of the approximated debiasing procedure in Algorithm 1. Algorithm 1 processes each data point only once and does not store any original raw data. In addition, the required summary statistics are at most $p$-dimensional. Therefore, Algorithm 1 is one-pass and only requests $\mathcal{O}(p)$ space complexity, as stated in Table 1.
Input : Streaming data $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, epoch schedule $\{T_{k}\}_{k\in[K]}$ and $\{T_{k}^{\prime}\}_{k\in[K^{\prime}]}$, regularization parameters $\{\lambda_{k}\}_{k\in[K]}$ and $\{\lambda_{k}^{\prime}\}_{k\in[K^{\prime}]}$, significance level $\alpha$, and index of variable of interest $j$.
@@
2 for $i=1,\ldots,n$ do
3      Receive the data point $(\bm{x}_{i},y_{i})$;
@@ 4       Conduct the RADAR with the hyper-parameters $\{T_{k},\lambda_{k}\}_{k\in[K]}$ ;
6       Obtain the online estimator $\widehat{\bm{\beta}}^{(i)}$ and ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}$ via (4) and (7) respectively;
7       Update $\bm{a}_{1}^{(i)},\bm{a}_{2}^{(i)},a_{3}^{(i)},a_{4}^{(i)},a_{5}^{(i)}$ recursively as (14);
8       Compute ${\widehat{\bm{\beta}}}^{(i)}_{j,\text{de}}$ and $\widehat{\nu}^{(i)}_{j}$ via (13) and (15) respectively;
@@ 9       Output the $(1-\alpha)$-confidence interval
10                 $\left({\widehat{\bm{\beta}}}^{(i)}_{j,\text{de}}-z_{\alpha/2}\widehat{\nu}^{(i)}_{j},{\widehat{\bm{\beta}}}^{(i)}_{j,\text{de}}+z_{\alpha/2}\widehat{\nu}^{(i)}_{j}\right)$ ;
11       Update $\left(\widehat{\bm{\beta}}^{(i)},{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)},\bm{a}_{1}^{(i)},\bm{a}_{2}^{(i)},a_{3}^{(i)},a_{4}^{(i)},a_{5}^{(i)}\right)$ and clear the raw data in memory;
@@
Algorithm 1 Pseudocode of the proposed ADL algorithm.
@@ 3.1 Oracle inequalities
To show the consistency of the RADAR output $\widehat{\bm{\beta}}^{(n)}$, we need the following assumptions.
@@
The response $Y$ is a sub-Gaussian random variable with the sub-Gaussian norm $\kappa_{2}$. In addition, the second derivative of the link function is Lipschitz and uniformly bounded by a constant $M_{\Phi}$, that is, $\max_{t\in\mathbb{R}}\lvert\ddot{\Phi}(t)\lvert\leq M_{\Phi}$.
@@
Assume $\lVert\bm{\beta}^{*}\lVert_{0}=s_{0}$ and $\lVert\bm{\beta}^{*}\lVert_{2}=m$ for a universal constant $m$.
Assumption 1 considers the sub-Gaussian random design, which is commonly adopted in establishing the oracle inequality for the high-dimensional model in the literature. Assumption 2 is a mild regularity condition on the generalized linear model (1). It is satisfied by a broad class of models, such as linear model with sub-Gaussian error, logistic model, and probit model. Assumption 3 is a basic assumption on the parameter space in the sparse regression settings.
@@
Suppose that Assumptions 1, 2 and 3 hold. With the choice of hyper-parameters in (20), the following events
@@
There exists a universal constant $M_{2}$ such that $0<M_{2}^{-1}\leq\Lambda_{\min}(\overline{\bm{\Theta}}(\bm{\beta}^{*}))\leq\Lambda_{\max}(\overline{\bm{\Theta}}(\bm{\beta}^{*}))\leq M_{2}<\infty$. The columns of $\overline{\bm{\Theta}}(\bm{\beta}^{*})$ is $s_{1}$-sparse: $\max_{j\in[p]}\lVert\{\overline{\bm{\Theta}}(\bm{\beta}^{*})\}_{\cdot,j}\lVert_{0}$  $=s_{1}.$
@@
There is a universal constant $L$ such that $\lVert\overline{\bm{\Theta}}(\bm{b}_{1})-\overline{\bm{\Theta}}(\bm{b}_{2})\lVert_{1}\leq L\lVert\bm{b}_{1}-\bm{b}_{2}\lVert_{1}$, for any $\bm{b}_{1},\bm{b}_{2}\in\{\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}^{*}\lVert_{1}\leq 2d_{0}.\}$
@@
In contrast to Theorem 3.1, the error bound in (18) not only involves the initial error $d_{1}$ and the sparsity $s_{1}$ of the inverse Hessian matrix, but also the error from the online lasso algorithm via $d_{0}$ and $s_{0}$. This phenomenon shows the difference between the RADAR algorithm and the adaptive RADAR algorithm. As shown in (8), the loss function in the adaptive RADAR algorithm is dynamically changing because of the most recent plug-in estimates from the online lasso procedure. In other words, the target values are instantly changing. Therefore, additional error will be introduced from the drifted target values. With careful design of the hyperparameters as in (22), we are able to control this additional error. Consequently, the overall error will converge at the rate of $\mathcal{O}(\sqrt{n})$.
@@ 3.2 Asymptotic normality
By Theorems 3.1 and 3.2, we can establish the asymptotic normality of ${\widehat{\bm{\beta}}}^{(k)}_{j,\text{de}}$ defined in (13).
@@
Compared to existing works in high-dimensional inferences, our theoretical results are different in two aspects. On one hand, the conditions in Theorem 3.3 are made at the population level. In contrast, most of the existing inference methods (van de Geer et al.,, 2014; Ning and Liu,, 2017; Ma et al.,, 2021; Shi et al.,, 2020; Luo et al., 2023a, ) rely on the bounded individual probability condition. For example, in logistic regression, the bounded individual probability condition takes the form that $\mathbb{P}(y_{i}=1|\bm{x}_{i})\in(c,1-c)$ for $i\in[n]$ and some positive constant $c$. However, such a condition can be easily violated if $\bm{x}_{i}$ follows a normal distribution with a slightly large variance. To the best of our knowledge, except for Guo et al., (2021) and Cai et al., (2023), most existing methods require the bounded individual probability condition. Without the bounded individual probability condition, our method is shown to provide valid inference under the Gaussian design. The removal of such a condition is not only a technical advancement, but also a methodological improvement as our method is able to handle more general cases.
On the other hand, we require slightly stronger conditions in terms of the sparsity levels $s_{0}$ and $s_{1}$, sample size $n$, and the dimension of features $p$. This is the trade-off between online and offline algorithms. Specifically, the online estimator may not be the optimiser of some loss functions. Hence, some optimality conditions, such as the Karush–Kuhn–Tucker (KKT) condition, cannot be used. Therefore, we need to impose slightly stronger conditions compared to its offline counterpart so as to establish the asymptotic properties; see Theorem 4.11 in Deshpande et al., (2023) and Theorem 1 in Han et al., (2024).
@@ 4 Numerical studies
In this section, we conduct simulation studies to examine the finite-sample performance of the ADL in high-dimensional logistic regression. Besides that, we provide a real data example to show the ability of our method in dealing with ultra-high-dimensional online data. All experiments are performed with an 11th Gen Intel Core i7-11370H processor and 16GB memory storage.
@@ 4.1 Simulated data
The performance is based on four criteria: (1) Absolute bias: the absolute distance between estimator and ground truth. (2) Converge probability of 95% confidence interval (3) Length of 95% confidence interval and (4) the computation times. Note that there are three different categories among $\bm{\beta}^{*}_{i},i\in[p]$, that is $0,1$ and $-1$. We choose three elements from each category and calculate the corresponding debiased estimator respectively. The averaged performances for each category are recorded. The experiment is conducted with 200 replications and the results are summarized in Tables 2 and 3.
@@ 4.2 Real data example: Identifying malicious URLs
Different from the prediction error, we are more interested in which attributes are significant to make the malicious URL distinguished from the benign one. We take one further step to conduct inference with the approximated debiased lasso algorithm. We evaluate our algorithm in two aspects: accuracy and computational efficiency. However, it is pretty hard to judge accuracy since all features are anonymized. For this reason, we add one fake feature, which is randomly generated from the standard normal distribution, as a benchmark. In addition to the fake feature, we also conduct inference for features 16 and 32. The results of the first-day data (with the size 20,000) are presented in Figure 4.
@@
In terms of computational efficiency, the online algorithm should be able to give the response for the current data prior to the arrival of the next data. The data in Ma et al., (2009) is collected in an online fashion, with approximately $4.32s$ for each. In contrast, the ADL costs $0.3s$ for processing each single data. This result reveals the efficiency of the ADL and its ability to conduct online inference with ultra-high-dimensional data.
@@ 5 Conclusion
High-dimensional inference has been well investigated in the last decade. However, it still lacks an efficient method to conduct the inference in a real online manner. In this work, we contribute to the literature by proposing the ADL to deal with the high-dimensional generalized linear model. While maintaining comparable accuracy to the offline method, the ADL requires much less time and space complexities.
Several directions deserve further investigation. First, the ADL is designed for an objective function with a lasso penalty. It is unclear what if some other penalty functions, such as SCAD (Fan and Li,, 2001) and MCP (Zhang,, 2010), are used. Second, our work focuses on the parametric models. It is interesting to investigate whether our approach can be extended to the model with a non-parametric component to conduct semi-parametric inference. Last but not least, this work assumes that sequentially collected data points are independent and identically distributed (i.i.d). How to deal with non-i.i.d. design, such as the covariate shift, under an online setting is still an open question.
Appendix A Proof related to the Convergence
A.1 Hyper-parameters used in proof
We first list the hyper-parameters. The constants $c_{i},i\in[4]$ are universal constants, whose exact value will be given in the supplementary material. Recall that $d_{0}=\lVert\bm{\beta}^{(0)}-\bm{\beta}^{*}\lVert_{1}$. Let $\omega_{1}=9\sqrt{\log\log p}$ and $R_{1}=c_{1}d_{0}$. Since we always take the zero point as the initial value, $R_{1}=c_{1}\lVert\bm{\beta}^{*}\lVert_{1}$. Moreover,
$R_{k}=R_{1}/(\sqrt{2})^{k-1},\ \xi_{k}=\kappa_{1}^{2}M_{\Phi}R_{k},\ \zeta_{k}^{2}=2\{\kappa_{2}\log(p)+M_{\Phi}^{2}R_{k}^{2}\log^{2}(p)\},$ (20)
and $\omega_{k}^{2}=\omega_{1}^{2}+24\log(k)$ for $k\in[K]$. Based on these hyper-parameters, we further define $T_{0}=1$ and
@@ A.1 Hyper-parameters used in proof
A.2 Lemmas
Let $\bar{F}(\bm{\beta};\lambda_{k}):=\mathbb{E}\{F_{n}(\bm{\beta};\lambda_{k})\}$ and
$\bm{\beta}^{(n_{k})}=\operatorname*{arg\,min\;}_{\bm{\beta}\in\mathbb{R}^{n}}\bar{F}(\bm{\beta};\lambda_{k}).$ (21)
We first show the convergence rate within a single epoch by the following lemma:
@@ A.2 Lemmas
Suppose that Assumptions 1,2,3 hold and $\lVert\widehat{\bm{\beta}}^{(n_{k-1})}-{\bm{\beta}}^{(n_{k})}\lVert_{1}\leq R_{k}.$ Then,
@@
Note that we choose $\widehat{\bm{\beta}}^{(0)}$ as the zero point. Since $\lVert\widehat{\bm{\beta}}^{(0)}-{\bm{\beta}}^{*}\lVert_{1}\leq R_{1}$ and $\lVert{\bm{\beta}}^{*}\lVert_{1}\geq\lVert{\bm{\beta}}^{(n_{1})}\lVert_{1}$, we obtain $\lVert\widehat{\bm{\beta}}^{(0)}-{\bm{\beta}}^{(n_{1})}\lVert_{1}\leq R_{1}$. Then, according to Lemma A.1, we obtain that
@@
Appendix B Convergence rate in nodewise lasso
Recall that $d_{1}=\lVert\widehat{\bm{\gamma}}_{\cdot,j}^{(0)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}.$ Let $\omega_{1}^{\prime}=9\sqrt{\log\log p}$, $c_{i}^{\prime},i\in[4]$ denote some universal constants, and
@@ Appendix B Convergence rate in nodewise lasso
$R_{k}^{\prime}=R_{1}^{\prime}/(\sqrt{2})^{k-1},\ \xi_{k}^{\prime}=\kappa_{1}^{2}M_{\Phi}R_{k}^{\prime},\ (\zeta_{k}^{\prime})^{2}=2\{\kappa_{1}M_{2}^{2}\log(p)+M_{\Phi}^{2}R_{k}^{\prime 2}\log^{2}(p)\},$ (22)
@@
We also provide the convergence rate of the nodewise lasso using the adaptive RADAR.
@@
Suppose that Assumptions 1–4 and (17) hold. If $\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k-1}^{\prime})}-\bm{\gamma}_{\cdot,j}^{(n_{k}^{\prime})}\lVert_{1}\leq R_{k}^{\prime}$, then
$\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k}^{\prime})}-\bm{\gamma}_{\cdot,j}^{(n_{k}^{\prime})}\lVert_{1}\leq R_{k+1}^{\prime},\ \lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k}^{\prime})}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}\leq R_{k+1}^{\prime}$ (23)
holds with probability at least $1-4\exp(-\omega_{k}^{\prime 2}/12)$.
@@
$\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(0)}-\bm{\gamma}_{\cdot,j}^{(n_{1}^{\prime})}\lVert_{1}$ $\leq\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(0)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}+\lVert\bm{\gamma}_{\cdot,j}^{*}-\bm{\gamma}_{\cdot,j}^{(n_{1}^{\prime})}\lVert_{1}\leq\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(0)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}+CR_{1}/\sqrt{2}\leq R_{1}^{\prime}.$
@@
$\lVert\widehat{\bm{\gamma}}_{\cdot,j}^{(i)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}\leq c_{7}s_{1}R_{1}^{\prime}\sqrt{\frac{(\log p)^{3}}{i-n_{1}}},i=n_{1}^{\prime},n_{1}^{\prime}+1,\ldots,n,$
@@
$\lVert\widehat{\bm{\gamma}}_{\cdot,j}^{(i)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}\leq c_{7}s_{1}R_{1}^{\prime}\sqrt{\frac{(\log p)^{3}}{i-n_{1}}},i=n_{1}^{\prime},n_{1}^{\prime}+1,\ldots,n,$
@@
We first state some facts about the length of the epoch, that is,
[…]

### v2

@@ 2.4 Variance and confidence interval
To construct a confidence interval, we need to further find the variance of ${\widehat{\bm{\beta}}}^{(m)}_{j,\text{de}}$. To this end, we design another scalar-valued summary statistic
$a_{5}^{(m)}=\sum_{i=n_{l}+1}^{m}(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})-y_{i}\right\}^{2}.$
Then the estimated standard error of the ADL estimator is computed as
$\widehat{\nu}^{(m)}_{j}=\sqrt{a_{5}^{(m)}}/|a_{4}^{(m)}|.$ (15)
Consequently, based on (13) and (15), the $(1-\alpha)$-confidence interval for $\bm{\beta}_{j}^{*}$ with $10<\alpha<$ 1 at time $m$ is given by:
$\left({\widehat{\bm{\beta}}}^{(m)}_{j,\text{de}}-z_{\alpha/2}\widehat{\nu}^{(m)}_{j},{\widehat{\bm{\beta}}}^{(m)}_{j,\text{de}}+z_{\alpha/2}\widehat{\nu}^{(m)}_{j}\right),\ m>n_{l},$ (16)
where $z_{\alpha/2}$ is the $\{1-(\alpha/2)\}$-quantile of the standard normal distribution. We present the pseudocode of the ADL in Algorithm 1. In numerical studies, we take the zero point as the initial value for both $\widehat{\bm{\beta}}^{(0)}$ and ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(0)}$. Algorithm 1 processes each data point only once and does not store any original raw data. In addition, the required summary statistics are at most $p$-dimensional. Therefore, Algorithm 1 is one-pass and only requests $\mathcal{O}(p)$ space complexity, as stated in Table 1.
Input : Streaming data $\{(\bm{x}_{m},y_{m})\}_{m=1}^{n}$, epoch schedule $\{T_{k}\}_{k\in[K]}$ and $\{T_{k}^{\prime}\}_{k\in[K^{\prime}]}$, regularization parameters $\{\lambda_{k}\}_{k\in[K]}$ and $\{\lambda_{k}^{\prime}\}_{k\in[K^{\prime}]}$, significance level $\alpha$, and index of variable of interest $j$.
@@
2 for $m=1,\ldots,n$ do
3      Receive the data point $(\bm{x}_{m},y_{m})$;
@@ 4       Conduct the RADAR with the hyper-parameters $\{T_{k},\lambda_{k}\}_{k\in[K]}$ ;
6       Obtain the online estimator $\widehat{\bm{\beta}}^{(m)}$ and ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(m)}$ via (4) and (7) respectively;
7       Update $\bm{a}_{1}^{(m)},\bm{a}_{2}^{(m)},a_{3}^{(m)},a_{4}^{(m)},a_{5}^{(m)}$ recursively as (14);
8       Compute ${\widehat{\bm{\beta}}}^{(m)}_{j,\text{de}}$ and $\widehat{\nu}^{(m)}_{j}$ via (13) and (15) respectively;
@@ 9       Output the $(1-\alpha)$-confidence interval
10                 $\left({\widehat{\bm{\beta}}}^{(m)}_{j,\text{de}}-z_{\alpha/2}\widehat{\nu}^{(m)}_{j},{\widehat{\bm{\beta}}}^{(m)}_{j,\text{de}}+z_{\alpha/2}\widehat{\nu}^{(m)}_{j}\right)$ ;
11       Update $\widehat{\bm{\beta}}^{(m)},{\widehat{\bm{\gamma}}_{\cdot,j}}^{(m)},\bm{a}_{1}^{(m)},\bm{a}_{2}^{(m)},a_{3}^{(m)},a_{4}^{(m)},a_{5}^{(m)}$ and clear the raw data in memory;
@@
Algorithm 1 Pseudocode of the proposed ADL.
@@ 3.1 Oracle inequalities
To show the consistency of the output $\widehat{\bm{\beta}}^{(n)}$ from the RADAR algorithm, we need the following assumptions.
@@
The response $Y$ is a sub-Gaussian random variable with the sub-Gaussian norm $\kappa_{2}$. In addition, the second derivative of the link function is Lipschitz and uniformly bounded by a universal constant $M_{2}$, that is, $\max_{t\in\mathbb{R}}\lvert\ddot{\Phi}(t)\lvert\leq M_{2}$.
@@
Assume $\lVert\bm{\beta}^{*}\lVert_{0}=s_{0}$ and $\lVert\bm{\beta}^{*}\lVert_{2}=M_{3}$ for a universal constant $M_{3}$.
Assumption 1 considers the sub-Gaussian random design, which is commonly adopted in establishing the oracle inequality for the high-dimensional model in the literature. Assumption 2 is a mild regularity condition on the generalized linear models (1). It is satisfied by a broad class of models, such as linear model with sub-Gaussian error, logistic model, and probit model. Assumption 3 is a basic assumption on the parameter space in the sparse regression settings.
@@
Suppose that Assumptions 13 hold. With the choice of hyper-parameters in (20), the following events
@@
There exists a universal constant $M_{4}$ such that $0<M_{4}^{-1}\leq\Lambda_{\min}(\overline{\bm{\Theta}}(\bm{\beta}^{*}))\leq\Lambda_{\max}(\overline{\bm{\Theta}}(\bm{\beta}^{*}))\leq M_{4}<\infty$. The columns of $\overline{\bm{\Theta}}(\bm{\beta}^{*})$ is $s_{1}$-sparse: $\max_{j\in[p]}\lVert\{\overline{\bm{\Theta}}(\bm{\beta}^{*})\}_{\cdot,j}\lVert_{0}$  $=s_{1}.$
@@
There is a universal constant $M_{5}$ such that $\lVert\overline{\bm{\Theta}}(\bm{b}_{1})-\overline{\bm{\Theta}}(\bm{b}_{2})\lVert_{1}\leq M_{5}\lVert\bm{b}_{1}-\bm{b}_{2}\lVert_{1}$, for any $\bm{b}_{1},\bm{b}_{2}\in\{\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}^{*}\lVert_{1}\leq 2d_{0}.\}$
@@
In contrast to Theorem 3.1, the error bound in (18) not only involves the initial error $d_{1}$ and the sparsity $s_{1}$ of the inverse Hessian matrix, but also the error from the online lasso algorithm via $d_{0}$ and $s_{0}$. This phenomenon shows the difference between the RADAR algorithm and the adaptive RADAR algorithm. As shown in (8), the loss function in the adaptive RADAR algorithm is dynamically changing because of the most recent plug-in estimates from the online lasso procedure. In other words, the target values are constantly changing. Therefore, an additional error will be introduced from the drifted target values. With careful design of the hyperparameters as in (21), we are able to control this additional error. Consequently, the overall error will converge at the rate of $\mathcal{O}(1/\sqrt{n})$.
@@ 3.2 Asymptotic normality
By Theorems 3.1 and 3.2, we can establish the asymptotic normality of ${\widehat{\bm{\beta}}}^{(n)}_{j,\text{de}}$ defined in (13).
@@
Compared to existing works in high-dimensional inferences, our theoretical results are different in two aspects. On one hand, the conditions in Theorem 3.3 are made at the population level. In contrast, most of the existing inference methods (van de Geer et al.,, 2014; Ning and Liu,, 2017; Ma et al.,, 2021; Shi et al.,, 2021; Luo et al., 2023a, ) rely on the bounded individual probability condition. For example, in logistic regression, the bounded individual probability condition takes the form that $\mathbb{P}(y_{i}=1|\bm{x}_{i})\in(c,1-c)$ for $i\in[n]$ and some positive universal constant $c$. However, such a condition can be easily violated if $\bm{x}_{i}$ follows a normal distribution with a slightly large variance. To the best of our knowledge, except for Guo et al., (2021) and Cai et al., (2023), most existing methods require the bounded individual probability condition. Without the bounded individual probability condition, our method is shown to provide valid inference under the Gaussian design. The removal of such a condition is not only a technical advancement, but also a methodological improvement as our method is able to handle more general cases.
On the other hand, we require slightly stronger conditions in terms of the sparsity levels $s_{0}$ and $s_{1}$, sample size $n$, and the dimension of features $p$. This is the trade-off between online and offline algorithms. Specifically, the online estimator may not be the optimizer of some loss functions. Hence, some optimality conditions, such as the Karush–Kuhn–Tucker (KKT) condition, cannot be used. Therefore, we need to impose slightly stronger conditions compared to its offline counterpart so as to establish the asymptotic properties; see Theorem 4.10 in Deshpande et al., (2023) and Theorem 1 in Han et al., (2024).
@@ 4 Numerical studies
In this section, we conduct simulation studies to examine the finite-sample performance of the ADL in high-dimensional logistic regression. Besides that, we provide a real data example to show the ability of our method in dealing with ultra-high dimensional online data. All experiments are performed with an 11th Gen Intel Core i7-11370H processor and 16GB memory storage.
@@ 4.1 Simulated data
The performance is based on four criteria: (1) Absolute bias: the absolute distance between estimator and ground truth. (2) Converge probability of 95% confidence interval (3) Length of 95% confidence interval and (4) the computation times. There are three different categories among $\bm{\beta}^{*}_{i},i\in[p]$, that is $0,1$ and $-1$. We choose three elements from each category and calculate the corresponding debiased estimator respectively. The averaged performances for each category are recorded. The experiment is conducted with 200 replications and the results are summarized in Tables 2 and 3.
@@ 4.2 Real data example: Identifying malicious URLs
Different from the prediction error, we are more interested in which attributes are significant to make the malicious URL distinguished from the benign one. We take one further step to conduct inference with the ADL method. We evaluate our algorithm in two aspects: accuracy and computational efficiency. However, it is pretty hard to judge accuracy since all features are anonymized. For this reason, we add one fake feature, which is randomly generated from the standard normal distribution, as a benchmark. In addition to the fake feature, we also conduct inference for features 16 and 32. The results of the first-day data (with the size 20,000) are presented in Figure 4.
@@
In terms of computational efficiency, the online algorithm should be able to give the response for the current data prior to the arrival of the next data. The data in Ma et al., (2009) is collected in an online fashion, with approximately $4.32s$ for each. In contrast, the ADL costs $0.3s$ for processing each single data. This result reveals the efficiency of the ADL and its ability to conduct online inference with ultra-high dimensional data.
@@ 5 Conclusion
High-dimensional inference has been well investigated in the last decade. However, it still lacks an efficient method to conduct the inference in a real online manner. In this work, we contribute to the literature by proposing the ADL to deal with the high-dimensional generalized linear models. While maintaining comparable accuracy to the offline method, the ADL requires much less time and space complexities.
Several directions deserve further investigation. First, the ADL is designed for an objective function with a lasso penalty. It is unclear what if some other penalty functions, such as SCAD (Fan and Li,, 2001) and MCP (Zhang,, 2010), are used. Second, our work focuses on the parametric models. It is interesting to investigate whether our approach can be extended to the model with a non-parametric component (Yang and Yao,, 2023; Quan and Lin,, 2024) to conduct semi-parametric inference. Last but not least, this work assumes that sequentially collected data points are independent and identically distributed (i.i.d). How to deal with non-i.i.d. design, such as the covariate shift, under an online setting is still an open question.
Appendix A Proof of Theorem 3.1
We first list the hyper-parameters used in the RADAR algorithm. The constants $c_{i},i\in[4]$ are universal constants. Recall that $d_{0}=\lVert\widehat{\bm{\beta}}^{(0)}-\bm{\beta}^{*}\lVert_{1}$. Let $R_{1}=c_{1}d_{0}$,
$R_{k}=R_{1}/(\sqrt{2})^{k-1},\ \xi_{k}=\kappa_{1}^{2}M_{2}R_{k},\ \zeta_{k}^{2}=2\{\kappa_{2}\log(p)+M_{2}^{2}R_{k}^{2}\log^{2}(p)\},$ (20)
$\omega_{1}=9\sqrt{\log\log p}$ and $\omega_{k}^{2}=\omega_{1}^{2}+24\log(k)$ for $k\in[K]$. Based on these hyper-parameters, we further define the length of each epoch: $T_{0}=1$ and
@@
Let $\bar{F}(\bm{\beta};\lambda_{k}):=\mathbb{E}\{F_{n}(\bm{\beta};\lambda_{k})\}$ and $\bm{\beta}^{(n_{k})}=\operatorname*{arg\,min\;}_{\bm{\beta}\in\mathbb{R}^{n}}\bar{F}(\bm{\beta};\lambda_{k}).$ Before proving Theorem 3.1, we present the convergence rate within a single epoch by the following lemma:
@@
Suppose that Assumptions 13 hold and $\lVert\widehat{\bm{\beta}}^{(n_{k-1})}-{\bm{\beta}}^{(n_{k})}\lVert_{1}\leq R_{k}.$ Then,
@@
The proof is based on the mathematical induction. Without loss of generality, we consider $\widehat{\bm{\beta}}^{(0)}$ is the zero point. Due to $\lVert\widehat{\bm{\beta}}^{(0)}-{\bm{\beta}}^{*}\lVert_{1}\leq R_{1}$ and $\lVert{\bm{\beta}}^{*}\lVert_{1}\geq\lVert{\bm{\beta}}^{(n_{1})}\lVert_{1}$, we obtain $\lVert\widehat{\bm{\beta}}^{(0)}-{\bm{\beta}}^{(n_{1})}\lVert_{1}\leq R_{1}$. Otherwise, with the choice of $\lambda_{1}$, one can still show $\lVert\widehat{\bm{\beta}}^{(0)}-{\bm{\beta}}^{(n_{1})}\lVert_{1}\leq R_{1}$.
Then, according to Lemma A.1,
@@
Appendix B Proof of Theorem 3.2
We list the hyper-parameters used in the adaptive RADAR algorithm. Recall that $d_{1}=\lVert\widehat{\bm{\gamma}}_{\cdot,j}^{(0)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}.$ Let $\omega_{1}^{\prime}=9\sqrt{\log\log p}$, $c_{i}^{\prime},i\in[4]$ denote some universal constants, and
@@
$R_{k}^{\prime}=R_{1}^{\prime}/(\sqrt{2})^{k-1},\ \xi_{k}^{\prime}=\kappa_{1}^{2}M_{2}R_{k}^{\prime},\ (\zeta_{k}^{\prime})^{2}=2\{\kappa_{1}M_{4}^{2}\log(p)+M_{2}^{2}R_{k}^{\prime 2}\log^{2}(p)\},$ (21)
@@
We provide the convergence rate of the nodewise lasso by using the adaptive RADAR algorithm in a single epoch.
@@
Suppose that Assumptions 15 and (17) hold. If $\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k-1}^{\prime})}-\bm{\gamma}_{\cdot,j}^{(n_{k}^{\prime})}\lVert_{1}\leq R_{k}^{\prime}$, then $\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k}^{\prime})}-\bm{\gamma}_{\cdot,j}^{(n_{k}^{\prime})}\lVert_{1}\leq R_{k+1}^{\prime},\ \lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k}^{\prime})}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}\leq R_{k+1}^{\prime}$ holds with probability at least $1-4\exp(-\omega_{k}^{\prime 2}/12)$.
@@
$\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(0)}-\bm{\gamma}_{\cdot,j}^{(n_{1}^{\prime})}\lVert_{1}$ $\leq\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(0)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}+\lVert\bm{\gamma}_{\cdot,j}^{*}-\bm{\gamma}_{\cdot,j}^{(n_{1}^{\prime})}\lVert_{1}\leq\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(0)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}+(M_{5}R_{1}/\sqrt{2})\leq R_{1}^{\prime}.$
@@
$\lVert\widehat{\bm{\gamma}}_{\cdot,j}^{(i)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}\leq c_{7}^{\prime}s_{1}R_{1}^{\prime}\sqrt{\frac{(\log p)^{3}}{i-n_{1}}},i=n_{1}^{\prime},n_{1}^{\prime}+1,\ldots,n,$
@@
$\lVert\widehat{\bm{\gamma}}_{\cdot,j}^{(i)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}\leq c_{7}^{\prime}s_{1}R_{1}^{\prime}\sqrt{\frac{(\log p)^{3}}{i-n_{1}}},i=n_{1}^{\prime},n_{1}^{\prime}+1,\ldots,n,$